For the function y = x2 - 4x - 5, how do you write the function in the form of y = a(x - h)2 + k?

Answer by stanbon(75887)

For the function y = x^2 - 4x - 5, how do you write the function in the form of y = a(x - h)^2 + k?
---------
It's called completing the square.
Rewrite as
x^2-4x+?=y+5?
Complete the square on the left side and maintain the equality, as follow:
x^2-4x+4 = y+5+4
Factor to get:
(x-2)^2 = y+9
Solve for y to get:
y=(x-2)^2 -9
Cheers,
Stan H.